OpenStudy (anonymous):

Integration with u substitution. \[\int \frac{x}{\sqrt{x + 2}} dx\]

OpenStudy (anonymous):

\(u = x + 2\) \(du = dx\) So \(x = u - 2\), I suppose? \[\int \frac{u - 2}{\sqrt{u}} du\]